### v1

# Towards a complete classification of holographic entropy inequalities

Abstract
We propose a deterministic method to find all holographic entropy inequalities and prove the completeness of our method. We use a triality between holographic entropy inequalities, contraction maps and partial cubes. More specifically, the validity of a holographic entropy inequality is implied by the existence of a contraction map, which we prove to be equivalent to finding an isometric embedding of a contracted graph. Thus, by virtue of the completeness of the contraction map proof method, the problem of finding all holographic entropy inequalities is equivalent to the problem of finding all contraction maps, which we translate to a problem of finding all image graph partial cubes. We give an algorithmic solution to this problem and characterize the complexity of our method. We also demonstrate interesting by-products, most notably, a procedure to generate candidate quantum entropy inequalities.
1 Introduction
Entropy inequalities constrain quantum states non-trivially. The first such inequality discovered was subadditivity, which constrains the entropies of the quantum states of a bipartite system and its componentsAraki:1970ba ,
[…]
In addition to subadditivity, there are other inequalities that are obeyed by all quantum states. We list some of them below Araki:1970ba ; ssa1973 ; weakmono2003 ; Headrick_2007 ,
[…]
Notably, the above inequalities are the only unconditional linear inequalities known to be obeyed by all quantum states.
A holographic entropy inequality(HEI) is a non-trivial constraint on entanglement entropies of holographic states, a proper subset of all quantum states, at leading order. Holographic states are those quantum states in conformal field theory that are consistent with the existence of a semi-classical gravity dual, via the $AdS/CFT$ correspondenceMaldacena1997 . More precisely, the Ryu-TakayanagiRyu2006 formula, a formula that relates the entanglement entropy of a boundary subregion $A$ to a bulk geometric quantity, namely, the minimal surface $\gamma_{A}$ in the bulk homologous to $A$, i.e.,
$S(A)=\frac{\operatorname{area}\gamma_{A}}{4G_{N}},$ (5)
applies for such states at leading order.
The simplest holographic entropy inequality that is not obeyed by all quantum states is the monogamy of mutual information(MMI)MMI2013 involving three parties, given by
[…]
MMI (41) taken together with subadditivity (1) inequalities between subsystem pairs $AB$, $AC$ and $BC$ form the holographic entropy cone(HEC)Bao:2015bfa for three regions $\mathcal{C}_{3}$. The work of Bao:2015bfa formalised the concept of the HEC and characterized it as a rational polyhedral cone, whose facets are identified as the minimal set of tight [^1] A HEI is a facet of the HEC iff there exists a codimension-1 set of linearly independent holographic entropy vectors for which the HEI saturates. We will use the terms facet HEI and tight HEI interchangeably. inequalities for a given number of parties. In the same work, five new inequalities were discovered that describe the facets of the HEC for five regions $\mathcal{C}_{5}$. It was verified in HernandezCuenca:2019wgh that these five inequalities (upto symmetry) taken together with the lower-party inequalities completely characterize $\mathcal{C}_{5}$. We give two examples of five-party inequalities below,
$\begin{split}&S(ABC)+S(ABD)+S(ACE)+S(BCD)+S(BCE)\geq\\ &S(A)+S(BC)+S(BD)+S(CE)+S(ABCD)+S(ABCE),\end{split}$ (7)
$\begin{split}&S(AD)+S(BC)+S(ABE)+S(ACE)+S(ADE)+S(BDE)+S(CDE)\geq\\ &S(A)+S(B)+S(C)+S(D)+S(AE)+S(DE)+S(BCE)+S(ABDE)+S(ACDE).\end{split}$ (8)
Over the past few years, several interesting directions emerged in the quest for discovering higher party HEIs to characterize $\mathcal{C}_{n}(n\geq 6)$. The work of Hernandez-Cuenca:2023iqh discovered 1877 novel six-party inequalities, where the search is motivated by the observation that all known tight HEIs (except subadditivity [^2] The subadditivity inequality is only balanced; it is neither superbalanced, nor expressible as a sum over (conditional) tripartite information.) are superbalancedHe:2020xuo and can be written as a sum of tripartite and conditional tripartite information, conveniently expressible in the $I$-basisHe:2019repackaged . A parallel development was the construction of the holographic cone of average entropies(HCAE)Czech:2021rxe which motivated the discovery of two additional families of HEIs [^3] The first family of cyclic inequalities was discovered in Bao:2015bfa valid for any odd number of parties $p$. For example, inequality 7 is the $p=5$ case. exploiting entanglement wedge nesting relationsCzech:2023xed ; Bao:2024vmy . Such geometrization of HEIs provided excellent opportunities to study various qualitative features of (a family of) HEIs. Another method for studying HEIs involved the use of bit-threads Freedman_2016 ; headrick2022covariant ; cui2019 . Several other works appeared in the recent years to study various aspects of the HECHernandez-Cuenca:2022marginal ; Fadel:2021urx ; Akers:2021lms ; Czech:2024rco and beyond the HEC, such as using hypergraphsBao:2020zgx ; Bao:2020uku , topological link modelsBao:2021gzu and cycle flowsHe:2023rox , to name a few. [^4] In all of these works, the entanglement entropy considered are bipartite entanglement measures between a sub-region $X$ and its complement $\bar{X}$ calculated for various multipartite arrangements of subregion $X$. Some other works concerning multipartite entropies calculated using multipartite entanglement measures include Gadde:2022cqi ; Gadde:2023zzj ; Penington:2022dhr ; Harper:2024ker .
Despite these developments, the task of fully characterizing all possible HEIs has proven elusive. Two key challenges stand in the way of this task. First, there is a combinatorial explosion associated with generating candidate inequalities. Second, one must develop methods, also often with similar combinatorial challenges, in order to prove a given conjectured inequality. One method of proving a HEI is to find a corresponding contraction map (which we define in subsection 2.2). Until recently, the computational complexity of finding such maps was doubly exponential in the number of terms on the left hand side (LHS) using greedy approaches. The work of Bao:2024contraction_map  [^5] Another complementary approach to speed-up the contraction map method can be found in Li:2022jji . reduced this complexity significantly to a single exponential and demonstrated the completeness of the contraction map method.
In this paper, we take a step towards a complete classification of HEIs. However, we step aside from the conventional classification of HEIs into classes of HEC $\mathcal{C}_{n}$ of a fixed party number $n$. Instead, we classify HEIs into the classes $\mathcal{H}_{M}$ identified by the number of LHS terms (with unit coefficient) $M$. For example, the five party inequalities 7 and 8 both belong to $\mathcal{C}_{5}$ but they have different numbers of LHS terms. In our classification, they will correspond to two different classes, $\mathcal{H}_{5}$ and $\mathcal{H}_{7}$ respectively.
The organization of this paper is as follows. We outline the problem statement motivated by the holographic entropy cone program in 2.1. In section 2.2, we define the ‘proof by graph contraction’ method after reviewing the ‘proof by contraction’ method and introducing several key concepts, such as partial cubes and graph contraction maps, from graph theory. In section 2.3, we rephrase the problem statement in binary expressions and graph theory. Moreover, we prove the equivalence of the problem statements by showing that the existence of a contraction map is a necessary and sufficient condition for that of a graph contraction map. In section 2.4, we give a basic framework of the algorithm as a solution to the problem statement in graph theory and argue their completeness. More specifically, we show that all relevant graphs can be constructed from graph contraction of a hypercube graph. In section 3, we give an algorithm to read off HEIs from a given contraction map (derived from a graph), followed by a classification of HEIs. Lastly, we discuss our results, their implications, and the future directions in section 4.
2 The problem statement: Find all holographic entropy inequalities
In general, a $n$-party holographic entropy inequality (HEI) involving $n$ disjoint regions $[n]:=\{A_{1},\cdots,A_{n}\}$ (and a purifier $O$), can be written in a basis of subregion entropies, constructed by their proper power set $P(\{A_{1},\cdots,A_{n}\})\backslash\emptyset$ containing $12^{n}-$ 1 possible elements. For example, in the case of the three regions $\{A,B,C\}$, the entropy basis is lexicographically written as
$\{S_{A},S_{B},S_{C},S_{AB},S_{AC},S_{BC},S_{ABC}\}$ (9)
An inequality $\mathcal{Q}$ can be expressed as
$\mathcal{Q}=\sum_{i=1}^{2^{n}-1}a_{i}S_{X_{i}}\geq 0,$ (10)
where $a_{i}$ is an integer coefficient associated with an entropy term $S_{X_{i}}$ corresponding to subregion $X_{i}$. Here, $a_{i}=0$ implies that the corresponding subregion entropy $S_{X_{i}}$ is absent from the inequality, whereas a positive (negative) $c_{i}$ implies that $S_{X_{i}}$ is present in the LHS (RHS). We can rearrange 10 to include only non-zero coefficients, to get an inequality of the form
$\sum_{i=1}^{l}c_{i}S_{X_{i}}\geq\sum_{j=1}^{r}d_{j}S_{X_{j}},$ (11)
where $c_{i}$ and $d_{j}$ are positive integers. The number of non-zero coefficients on the LHS and RHS are $l$ and $r$ respectively. We would like to expand each coefficient, and re-write the inequality with repeating subregion entropies $S_{X_{i}}$ with unit coefficient, giving us
$\sum_{u=1}^{M}S_{L_{u}}\geq\sum_{v=1}^{N}S_{R_{v}}.$ (12)
The number of terms (after expanding the coefficients) are $M$ and $N$ respectively, i.e., $\sum_{i}c_{i}=M$ and $\sum_{j}d_{j}=N$.
2.1 Problem statement in physics
The thematic problem statement of our work is to find all holographic entropy inequalities.
Thematic Problem Find all holographic entropy inequalities.
We begin by fixing the number of LHS terms of our inequality (not the terms), i.e, we are fixing $M$. Since $M$ is arbitrary, finding all possible HEIs for a given $M$ allows us to find all possible HEIs. Now we are ready to define our main problem statement.
Main Problem 1 (Physics statement).
Consider a set of subregions. Given a convex composition of entropies with positive integer coefficients $c_{i}\geq 0$ for $\forall i=\{1,\cdots,l\}$,
$c_{1}S_{L_{1}}+\cdots+c_{l}S_{L_{l}}.$ (13)
Find all the possible convex combinations of entropies with positive integer coefficients [^6] It suffices to consider integer coefficients because the HEC is a rational polyhedral coneBao:2015bfa . $d_{j}\geq 0$ for $\forall j=\{1,\cdots,r\}$,
$d_{1}S_{R_{1}}+\cdots+d_{r}S_{R_{r}},$ (14)
such that
$c_{1}S_{L_{1}}+\cdots+c_{l}S_{L_{l}}\geq d_{1}S_{R_{1}}+\cdots+d_{r}S_{R_{r}}$ (15)
is a valid HEI.
We can simplify this problem a little bit to fit our methods by expanding the coefficients, such that the there are $M$ entropy terms on the LHS with unit coefficient. [^7] Note that when the coefficients $c_{i}>1$, there are repetition of terms. In fact, it suffices to stick to unit coefficients with non-repeating LHS terms, as all inequalities with $c_{i}>1$ can be uplifted from an $n$-party inequality to a $(n+k)$-party inequality with non-negative integer $k$ having non-repeating terms with unit coefficients on the LHS (see corollary 2.1). From now on, all inequalities we will refer to have unit coefficients. We thus reformulate our main problem 1 in the box 1. Note that this reformulation of the problem is equivalent to the original one.
Main Problem 1 (Unimodular)
Main Problem 1 (Physics statement: Unimodular).
Given $M$ LHS terms with unit coefficients, $S_{L_{1}}+\cdots+S_{L_{M}}.$ (16) Find all possible RHS terms, such that the following inequality holds $S_{L_{1}}+\cdots+S_{L_{M}}\geq S_{R_{1}}+\cdots+S_{R_{N}},$ (17) for some $N$ with unit coefficients.
We will give an algorithmic recipe to answer this problem in the language of graph theory. The LHS comprising $M$ terms (16) can be mapped to a hypercube graph $H_{M}$ of $2^{M}$ binary bitstrings. One then performs graph contractionsErciyes2018 on this hypercube graph $H_{M}$, resulting in a contracted graph is $G$. Since the number of possible graph contractions is finite (albeit, large), the number of contracted graphs is also finite. Isometric hypercube embeddability of this resultant graph $G$ into a hypercube $H_{N}$, should it exist, would define a contraction map from $\{0,1\}^{M}$ to $\{0,1\}^{N}$. Given a contraction map, one can find the HEIs corresponding to the contraction map by assigning boundary conditions. Thus, the problem of finding all inequalities can be framed as a problem of finding all contraction maps. We show the problem of finding all contraction maps is equivalent to determining which subset of graphs $G$ are partial cubes, where the graphs $G$ are obtained by graph contractions starting from $H_{M}$.
In the rest of this paper, we will prove that this method is sufficient to generate all possible HEIs. We will give a more formal mathematical presentation of the problem statements below after introducing relevant definitions. We summarize our algorithmic method in algorithm 1 in subsection 2.4.
2.2 Contraction maps and graph contraction maps
The ‘proof by contraction’ method was proposed in Bao:2015bfa to prove the validity of a candidate HEI (11). This method relies on the equivalence between holographic geometries and graphs, labelled by bitstrings. In principle, this method is based on the inclusion/exclusion of bulk regions and extracts the information of contributions of the RT surfaces to the convex compositions of the RT entropies and encodes them into bitstrings.
A particularly relevant set of bitstrings are those labeling regions adjacent to boundary subsystems; for each $i\in[n+1]$,
$x^{u}_{A_{i}}=\begin{cases}1,\qquad\text{if}\quad A_{i}\in L_{u}\\ 0,\qquad\text{otherwise,}\end{cases}$ (18)
where the ($n+1$)-th bitstring is all zeroes and is associated with the purifier $O$. These bitstrings are the boundary conditions (also known as occurrence bitstrings) and can be analogously defined for RHS subsystems using bitstrings $y\in\{0,1\}^{N}$.
A contraction map is a map between bitstrings from $\{0,1\}^{M}$ to $\{0,1\}^{N}$ such that the Hamming distance between every pair of bitstrings in $\{0,1\}^{M}$ is greater than or equal to the Hamming distance between their images in $\{0,1\}^{N}$. A contraction map learns about a particular inequality through the boundary conditions. We will now define these ideas more formally below.
Definition 2.1 (Hamming distance).
The Hamming distance between two bitstrings of length $L$ is defined as the number of positions where the bitstrings differ, i.e.,
$d_{H}(x,x^{\prime})=\sum_{u=1}^{M}|x^{u}-x^{\prime u}|.$ (19)
where $x^{u}$ is the $u$-th bit.
Theorem 2.1 (‘Proof by contraction’).
Bao:2015bfa Let $f:\{0,1\}^{M}\to\{0,1\}^{N}$ be a contraction map, i.e.,
$d_{H}(x,x^{\prime})\geq d_{H}(f(x),f(x^{\prime})),\;\forall x,x^{\prime}\in\{0,1\}^{M}.$ (20)
If $f(x_{A_{i}})=y_{A_{i}}$ for $\forall i\in\{1,\cdots,n+1\}$, then (12) is a valid $n$-party HEI.
We can rephrase the language of bitstrings into its natural habitat in graph theory by reinterpreting the bitstrings $\{0,1\}^{M}$ to be a $M$-dimensional hypercube, which is mapped to a subset of $\{0,1\}^{N}$ satisfying the contraction condition (20) to be a subgraph of $N$-dimensional hypercube.
Consider two unweighted and undirected graphs, $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$. In general, the graph map $\Phi:=(\phi^{V},\phi^{E})$ from $G_{1}$ to $G_{2}$ consists of a map $\phi^{V}:V_{1}\to V_{2}$ and $\phi^{E}:E_{1}\to E_{2}$. A weak graph homomorphism $\Phi$ is a graph map that either preserves the adjacency $(\phi^{V}(v),\phi^{V}(v^{\prime}))\in E_{2}$ or contracts the vertices $\phi^{V}(v)=\phi^{V}(v^{\prime})$ for $v,v^{\prime}\in V_{1}$ and $(v,v^{\prime})\in E_{1}$hammack2011 . Hence, it does not increase the graph distance. A graph distance $d(v,v^{\prime})$ is the shortest path between the vertices $v,v^{\prime}\in V$Bondy-Murty .
We define a graph contraction map to be a weak graph homomorphism to construct two alternative formulations below. Going forward, we will only use the term graph contraction map for clarity.
Definition 2.2 (A graph contraction map/weak graph homomorphism).
For a graph $G=(V,E)$, a graph map $\Phi:G\to\Phi(G)$ is a graph contraction map if
$d(v,v^{\prime})\geq d(\phi^{V}(v),\phi^{V}(v^{\prime})),\;\forall v,v^{\prime}\in V.$ (21)
From theorem 2.1, the domain of graph contraction maps is always an $M$-dimensional hypercube $H_{M}$. Moreover, by virtue of the bitstrings picture, the images $\Phi(H_{M})$ of the graph contraction maps $\Phi$ are subgraphs of the $N$-dimensional hypercube $H_{N}$. In fact, the subgraphs $\Phi(H_{M})$ should be isometrically embeddable in $H_{N}$ [^8] A graph $G$ is isometrically embeddable to another graph $G^{\prime}$ if the adjacency and the graph distance of all pairs of vertices of $G$ are preserved.. Such graphs are called partial cubes. We define a partial cube and its isometric dimension below.
Definition 2.3 (Partial cubeOvchinnikov ; WINKLER1984221 ).
A graph $G$ is a partial cube if $G$ is isometrically embeddable to a $D$-dimensional hypercube graph $H_{D}$.
Definition 2.4 (Isometric dimension Ovchinnikov ).
The isometric dimension $dim(G)$ of a partial cube $G$ is the minimum dimension of a cube in which $G$ is isometrically embeddable.
A $M$-dimensional hypercube graph is the trivial partial cube of isometric dimension $dim(H_{M})=M$. We denote $\Phi:H_{M}\to H_{N}$ if a graph map is from a $M$-dimensional hypercube $H_{M}$ to a partial cube with $dim(\Phi(H_{M}))=N$.
Correspondingly to the occurrence bitstrings(18), we define occurence vertices or boundary vertices as
$\big{(}\iota_{M}(v_{A_{i}})\big{)}^{u}=\begin{cases}1,\qquad\text{if}\quad A_{i}\in L_{u}\\ 0,\qquad\text{otherwise,}\end{cases}$ (22)
where $\iota_{M}:V_{M}\to\{0,1\}^{M}$ is an isometry, i.e.,
$d(v,v^{\prime})=d_{H}(\iota_{M}(v),\iota_{M}(v^{\prime}))$ (23)
for $v,v^{\prime}\in V_{M}$.
Now, we are ready to define the ‘proof by graph contraction’, as follows.
Theorem 2.2 (‘Proof by graph contraction’).
Bao:2015bfa Let $\Phi:H_{M}=(V_{M},E_{M})\to H_{N}=(V_{N},E_{N})$ be a graph contraction map, i.e.,
$d(v,v^{\prime})\geq d(\phi^{V}(v),\phi^{V}(v^{\prime})),\;\forall v,v^{\prime}\in V_{M}.$ (24)
If $\phi^{V}(v_{A_{i}})=v_{A_{i}}$ for $\forall i\in\{1,\cdots,n+1\}$, then (11) is a valid HEI.
2.3 Problem statements in mathematics: bitstrings and graph theory
In this subsection, we define and discuss two alternative formulations of the main problem 1 as problems 2 and 3 and prove their equivalence. In the language of bitstrings, the main problem 1 is rephrased as follows.
Main Problem 2 (Bitstrings)
Main Problem 2 (Mathematical statement: Bitstrings).
Given a complete set of bitstrings [^9] A complete set of bitstrings is the set of all possible $2^{M}$ binary bitstrings of length $M$. $\{0,1\}^{M}$ for a fixed $M\in\mathbb{Z}_{+}$, find the set of all the possible contraction maps =$\mathcal{F}=\{f|f:\{0,1\}^{M}\to\{0,1\}^{N}\}$ for some $N\in\mathbb{Z}_{+}$, where $f$ maps the complete set of bitstrings in $\{0,1\}^{M}$ to a subset of bitstrings in $\{0,1\}^{N}$.
The equivalence between the problem 1 and the problem 2 is guaranteed by theorem 2.1 and the completeness of ‘proof by contraction’Bao:2024contraction_map which states that there exists at least one contraction map for every valid HEI.
In the graph-theoretic reformulation, we have the following problem statement.
Main Problem 3 (Graph theory)
Main Problem 3 (Mathematical statement: Graph theory).
Given a $M$-dimensional hypercube $H_{M}$ for a fixed $M\in\mathbb{Z}_{+}$, find the set of all the possible graph contraction maps $\mathcal{F}_{G}=\{\Phi|\Phi:H_{M}\to H_{N}\}$ such that the images $Im\mathcal{F}_{G}=\{\Phi(H_{M})\}$ are partial cubes of $dim(\Phi(H_{M}))=N$ for some $N\in\mathbb{Z}_{+}$.
We prove that problem 2 is equivalent to the problem 3 by following proposition. Therefore, all the three main problems 1, 2, and 3 are equivalent.
Proposition 2.1 (Equivalence between contraction maps and graph contraction maps).
Given $M\in\mathbb{Z}_{+}$, there exists a contraction map $f:\{0,1\}^{M}\to\{0,1\}^{N}$ for some $N\in\mathbb{Z}_{+}$ if and only if there exists a graph contraction map $\Phi:H_{M}\to H_{N}$.
Proof.
$(f\rightarrow\Phi)$ We first show that if $f:\{0,1\}^{M}\to\{0,1\}^{N}$ is a contraction map given $M$ for some $N$, there exists a graph contraction map $\Phi:H_{M}=(V_{M},E_{M})\to H_{N}=(V_{N},E_{N})$.
Let us define an isometry $\tilde{\iota}_{J}:\{0,1\}^{J}\to V_{J}$ for $J\in\mathbb{Z}_{+}$ such that
$d_{H}(x,x^{\prime})=d(\tilde{\iota}_{J}(x),\tilde{\iota}_{J}(x^{\prime})),\forall x,x^{\prime}\in\{0,1\}^{J}.$ (25)
Then, we construct a graph map $\Phi_{f}:=(\phi^{V}_{f},\phi^{E}_{f})$ from a contraction map $f$ using the isometries $\tilde{\iota}_{M}:\{0,1\}^{M}\to V_{M}$ and $\tilde{\iota}_{N}:\{0,1\}^{N}\to V_{N}$. That is, we define $\phi^{V}_{f}$ and $\phi^{E}_{f}$ such that, for $x\in\{0,1\}^{M}$ and $f(x)\in\{0,1\}^{N}$,
  • $\phi^{V}_{f}\circ\tilde{\iota}_{M}(x)=\tilde{\iota}_{N}\circ f(x)$
  • $\phi^{E}_{f}(\tilde{\iota}_{M}(x),\tilde{\iota}_{M}(x^{\prime}))=(\tilde{\iota}_{N}\circ f(x),\tilde{\iota}_{N}\circ f(x^{\prime}))$ if
    $d_{H}(f(x),f(x^{\prime}))=1,$ (26)
    however $\phi^{E}(E)/\phi^{E}_{f}(\tilde{\iota}_{M}(x),\tilde{\iota}_{M}(x^{\prime}))$ [^10] We denote $G/e$ as deleting an edge $e$ from a graph $G$. We have this condition to remove self-loops from the image graph $\Phi(H_{M})$. if
    $d_{H}(f(x),f(x^{\prime}))=0.$ (27)
By construction, $\phi^{V}_{f}\circ\tilde{\iota}_{M}(x)\in V_{N}$ and $\phi^{E}_{f}(\tilde{\iota}_{M}(x),\tilde{\iota}_{M}(x^{\prime}))\in E_{N}$. Thus, $\Phi_{f}$ is a graph map from $H_{M}$ to $H_{N}$. In other words, the image $\Phi_{f}(H_{M})$ is always a partial cube of $dim(\Phi_{f}(H_{M}))=N$.
Figure 1: The diagram of the maps in the proof of proposition 2.1. The isometries $\iota_{M,N}$ and $\tilde{\iota}_{M,N}$ map between bitstrings $\{0,1\}^{M,N}$ and the set of vertices $V_{M,N}$ of hypercubes $H_{M,N}$. $f$ is a contraction map. $\phi^{V}$ is a vertex map from $H_{M}$ to $H_{N}$.
$\Phi_{f}$ is a graph contraction map because, for all $x,x^{\prime}\in\{0,1\}^{M}$,
$\begin{split}d(\tilde{\iota}_{M}(x),\tilde{\iota}_{M}(x^{\prime}))&=d_{H}(x,x^{\prime})\\ &\geq d_{H}(f(x),f(x^{\prime}))\\ &=d(\phi^{V}_{f}\circ\tilde{\iota}_{M}(x),\phi^{V}_{f}\circ\tilde{\iota}_{M}(x^{\prime})).\\ \end{split}$ (28)
We applied the definition of isometry $\tilde{\iota}_{M}$ for the first equality. The second inequality is due to the contraction property of $f$. The isometry $\tilde{\iota}_{N}$ and the definition of $\phi^{V}_{f}$ implies the last equality. Therefore, $\Phi_{f}:H_{M}\to H_{N}$ is a graph contraction from $H_{M}$ to $H_{N}$.
$(f\leftarrow\Phi)$ We next show that if a graph map $\Phi$ is a graph contraction map $\Phi:H_{M}=(V_{M},E_{M})\to H_{N}=(\phi^{V}(V_{M}),\phi^{E}(E_{M}))$ given $M$ for some $N$, there exists a contraction map $f:\{0,1\}^{M}\to\{0,1\}^{N}$.
Similarly, we define an isometry $\iota_{J}:V_{J}\to\{0,1\}^{J}$ for $J\in\mathbb{Z}_{+}$ such that
$d_{H}(\iota_{J}(v),\iota_{J}(v^{\prime}))=d(v,v^{\prime}),\forall v,v^{\prime}\in V_{J}$ (29)
For $\iota_{M}:V_{M}\to\{0,1\}^{M}$ and $\iota_{N}:V_{N}\to\{0,1\}^{N}$, a contraction map $f_{\Phi}$ is constructed such that
$f_{\Phi}\circ\iota_{M}(v)=\iota_{N}\circ\phi^{V}(v).$ (30)
Then, we see that $f_{\Phi}:\{0,1\}^{M}\to\{0,1\}^{N}$ simply because $\iota_{M}(v)\in\{0,1\}^{M}$ and $\iota_{N}\circ\phi^{V}(v)\in\{0,1\}^{N}$. For all $v,v^{\prime}\in V_{M}$,
$\begin{split}d_{H}(\iota_{M}(v),\iota_{M}(v^{\prime}))&=d(v,v^{\prime})\\ &\geq d(\phi^{V}(v),\phi^{V}(v^{\prime}))\\ &=d_{H}(f_{\Phi}\circ\iota_{M}(v),f_{\Phi}\circ\iota_{M}(v^{\prime})).\\ \end{split}$ (31)
Therefore, $f_{\Phi}:\{0,1\}^{M}\to\{0,1\}^{N}$ is a contraction map.
2.4 An algorithmic solution to the main problems
In this subsection, we provide an algorithmic solution to the main problem 3. The key ingredients of our solution are i) graph contraction maps and ii) partial cubes or isometric hypercube embeddability. We use the following three algorithms to realize a graph contraction map from a $M$-dimensional hypercube graph $H_{M}=(V_{M},E_{M})$ to a partial cube $G$ of $dim(G)=N$, as follows: (i) partition generatorer1988 , (ii) graph contractionErciyes2018 , and (iii) partial cube identifierEppstein_2011 , see algorithm 1 and figure 2.
The Partition generator algorithm in er1988 recursively generates all partitions of a given set, a set $V_{M}$ of vertices of hypercube graph $H_{M}$ in our case. The number of partitions is given by Bell number $B_{\alpha}$ of a set of cardinality $\alpha$, i.e.,
$B_{\alpha}=\sum_{\beta=0}^{\alpha}P(\alpha,\beta)$ (32)
where $P(\alpha,\beta)$ is the Stirling number of the second kind computing the number of partitions of a set into $\beta$ non-empty subsets.
The Graph contraction algorithm Erciyes2018 has been studied to provide parallel algorithms to solve graph-associated problems, for instance, Erciyes2018 ; LOMBARDI2022429 ; Guattery1992 . We apply the algorithm to a hypercube graph $H_{M}$ to generate contracted graphs based on the partitions generated above.
However, it turns out that the output graphs generated by the Graph contraction algorithm are not necessarily isometrically hypercube embeddable or partial cubes. So, we apply a polynomial time algorithm Eppstein_2011 to check whether the graph is a partial cube. We call this algorithm Partial cube identifier.
We now describe algorithm 1 in detail. For a hypercube graph $H_{M}=(V_{M},E_{M})$ with a coloring map $\iota_{M}:V_{M}\to\{0,1\}^{M}$, let us call $(H_{M},\iota_{M})$ as an initial graph data.
  1. 1.
    Partition generatorer1988 : Given an initial graph data $(H_{M},\iota_{M})$, generate $B_{2^{M}-1}$ partitions, $p^{\sigma}$ for $\sigma=1,\cdots,B_{2^{M}-1}$, of $V_{M}$, i.e.,
    $\bigcup_{w}V_{p^{\sigma}_{w}}\cup\{0\}^{M}=V_{M}$ (33)
    where $V_{p^{\sigma}_{w}}$ are disjoint subsets of vertices for the $\sigma$-th partition $p^{\sigma}$. $w=1,\cdots,\beta_{\sigma}$ labels the $\beta_{\sigma}$ non-empty subsets in the $\sigma$-th partition. We fix $\{0\}^{M}$ as a single element subset [^11] In our gauge choice, we choose $\{0\}^{M}$. One can choose, for instance, $\{1\}^{M}$ instead..
  2. 2.
    Choose a partition of $V_{M}$, e.g., $\bigcup_{w}V_{p^{\sigma}_{w}}\cup\{0\}^{M}=V_{M}$.
  3. 3.
    Graph contraction Erciyes2018 : Construct a graph $G_{\sigma}=(V_{\sigma},E_{\sigma})$.
    • Each vertex subset $V_{p^{\sigma}_{w}}$ in the partition becomes a vertex of $G_{\sigma}$, i.e., $V_{p^{\sigma}_{w}}\in V_{\sigma}$.
    • $(V_{p^{\sigma}_{w}},V_{p^{\sigma}_{w^{\prime}}})\in E_{\sigma}$.
    • Remove edges $(v,v^{\prime})$ between $v,v^{\prime}\in V_{p^{\sigma}_{w}}$.
    • Merge multiple parallel edges between $V_{p^{\sigma}_{w}}$ and $V_{p^{\sigma}_{w^{\prime}}}$.
  4. 4.
    Partial cube identifierEppstein_2011 : Check if $G_{\sigma}$ is a partial cube.
    • If so, the algorithm determines the isometric dimension $dim(G_{\sigma})=N$ for some $N\in\mathbb{Z}_{+}$ and the coloring map $\iota_{N}:V_{\sigma}\to\{0,1\}^{N}$. Thus, we obtain a final graph data $(G_{\sigma},\iota_{N})$ depending on a choice of partition $\sigma$.
    • Otherwise, discard the non-partial cubes $G_{\sigma}$.
Figure 2: Reading off a contraction map from the graph contraction mapping from $H_{M=3}$ to $H_{N=4}$. The blue dots are vertices. The black lines are the edges connecting the vertices. (a) Each vertex is colored by $\{0,1\}^{3}$. A partition is chosen as $\{000\},\{011\},\{101\},\{110\},\{001,010,100,111\}$. The vertices colored with the bitstrings in $\{001,010,100,111\}$ are enclosed by a rounded square. (b) The graph contraction with the choice of partition generates a star graph, a partial cube of isometric dimension $4$. The vertices get a coloring with a subset of $\{0,1\}^{4}$, e.g., $\{0000,0001,0011,0101,1001\}$.
Once we generate all partial cubes with all possible coloring maps, we can read off a contraction map from the initial graph data $(H_{M},\iota_{M})$ and the final graph data $(G_{\sigma},\iota_{N})$. That is, for a choice of partition, a contraction map $f_{(\sigma,\iota_{N})}:\{0,1\}^{M}\to\{0,1\}^{N}$ is determined [^12] See, for example, figure 2. such that
$f_{(\sigma,\iota_{N})}\circ\iota_{M}(v)=\iota_{N}(V_{p_{\sigma}})\text{ for $\forall v\in V_{p_{\sigma}}$ and $\forall V_{p_{\sigma}}\in V_{\sigma}$}.$ (34)
Algorithm 1 A road map to generate all HEIs for a fixed number of LHS terms.
1:procedure All HEIs for fixed $M$
2:     Construct an initial graph data $(H_{M},\iota_{M})$.
3:     Generate all partitions $\{\sigma\}$ of hypercube vertices using Partition generator.
4:     for  $\sigma\in\{\sigma\}$  do
5:         Find the contracted graph $G_{\sigma}=(V_{\sigma},E_{\sigma})$ using Graph contraction.
6:         Obtain a final graph data $(G_{\sigma},\iota_{N})$ by checking whether $G_{\sigma}$ is a partial cube using Partial cube identifier.
7:         if $G_{\sigma}$ is a partial cube then
8:              Generate HEIs from $(G_{\sigma},\iota_{N})$ using Contraction Map to Inequalities.
9:         end if
10:     end for
11:end procedure
The procedure generates all possible contraction maps because, first, it generates all possible partial cubes from $H_{M}$. Moreover, the contraction maps depend on both a choice of partition and a coloring map. The graph data $(G_{\sigma},\iota_{N})$ and $(G_{\sigma},\iota^{\prime}_{N})$ that differ only by the coloring maps $\iota_{N}\neq\iota_{N}^{\prime}$ can generate possibly two distinct contraction maps. In addition, even if two partial cubes $G_{\sigma}$ and $G_{\sigma^{\prime}}$ of the graph data $(G_{\sigma},\iota_{N})$ and $(G_{\sigma^{\prime}},\iota_{N})$ for $\sigma\neq\sigma^{\prime}$ are graph isomorphic and the coloring maps $\iota_{N}$ are identical, they could also construct different contraction maps. Hence, the algorithm considers all possibilities of contraction maps.
However, the algorithmic solution above is overcomplete because it could generate identical contraction maps. Thus, the total complexity of the algorithms is overestimated, as discussed in section 4.
In section 3, we discuss how to read off inequalities by choosing the number of parties $n$ and their boundary conditions for a fixed number of terms on the LHS and RHS of inequalities, given a contraction map $f$ constructed by the above procedure. This allows us to generate valid $n$-party HEIs. Similarly, we can choose another $n^{\prime}$ and obtain $n^{\prime}>n$-party HEIs from the contraction map $f$. Starting from $n^{\prime}$-party inequality, we can always arrive at a $n$-party HEI, possibly with non-unit coefficients, by removing $n^{\prime}-n$ boundary conditions.
Since we can generate all contraction maps, and hence, all HEIs, there should exist at least one $n^{\prime}>n$-inequality that lands on a target $n$-party HEI by removing $n^{\prime}-n$ boundary conditions. Thus, we have the following corollary of this section.
Corollary 2.1.
All $n$-party HEIs with non-unit coefficients are generated from $n^{\prime}$-party HEIs with unit coefficients for some $n^{\prime}>n$.
We will illustrate this with an example. Consider the following five-party facet inequality,
$\begin{split}&2S(ABC)+S(ABD)+S(ABE)+S(ACD)+S(ADE)+S(BCE)+S(BDE)\geq\\ &S(AB)+S(AC)+S(AD)+S(BC)+S(BE)+S(DE)+S(ABCD)+S(ABCE)+S(ABDE)\end{split}$ (35)
It has a term $2S(ABC)$, which we split as
$2S(ABC)\rightarrow S(ABCF)+S(ABCG)$
Balancing the inequality on both sides, we can generate the following seven-party inequality,
$\begin{split}&S(ABCF)+S(ABCG)+S(ABDF)+S(ABEG)+S(ACD)+S(ADE)+S(BCE)+S(BDE)\geq\\ &S(AB)+S(AC)+S(AD)+S(BC)+S(BE)+S(DE)+S(ABCDF)+S(ABCEG)+S(ABDEFG)\end{split}$ (36)
which can be easily proved using Bao:2024contraction_map .
3 Holographic entropy inequalities from contraction maps
3.1 A greedy algorithm to generate HEIs from a contraction map
In this section, we describe how to construct HEIs from a given contraction map. A contraction map is a map between $H_{M}$ and $H_{N}$, and a priori, such a map doesn’t know about the inequality. As mentioned earlier, the knowledge of the inequality is imparted to the contraction map by virtue of the boundary conditions(18), where one assigns $(n+1)$ bitstrings to $n$ monochromatic regions (and the all $\{0\}^{M}$ bitstring is assigned to the purifier $O$).
In principle, for an inequality involving $n$-parties (and a purifier) with $M$ terms on the LHS, the number of combinations in which one can assign the boundary conditions is $C(2^{M}-1,n)$ [^13] We do not get any new HEI by permutations, only the labels of regions are exchanged.. Each such choice gives us a HEI (not necessarily unique). We are interested in balanced [^14] One may also impose the condition of superbalance to narrow down the search for facet inequalities. If we are interested only in true but not necessarily facet inequalities, superbalance is not required. inequalities that have $M(N)$ non-trivial columns on the LHS(RHS). So we add the criterion that after the assignment of occurrence bitstrings, all columns must correspond to some non-trivial subregion entropy. We can then store all the unique HEI candidates. They are all valid HEIs by the completeness of contraction map. One may further check using known extreme rays if these HEIs are potentially facets of the HEC. However, in our present work, we are not interested in that step. We summarize our algorithm to read off HEIs from the contraction map below in algorithm 2 and give an example of this exercise in 3.3. The computational complexity of this procedure goes as $\mathcal{O}\left((M+N)\left(2^{M}\right)^{n}\right)$ when $n<<2^{M}$.
Algorithm 2 A greedy algorithm to generate HEIs from a contraction map.
1:procedure Contraction Map to Inequalities
2:     Read the contraction map $H_{M}\rightarrow H_{N}$.
3:     Assign the $\{O\}^{M}$ bitstring to O.
4:     For $n$-parties, generate $\texttt{num}=C(2^{M}-1,n)$ combinations of boundary conditions.
5:     for $i=0;\ i<\texttt{num};\ i++$ do
6:         Use boundary conditions to generate HEI $\mathcal{Q}_{i}$.
7:         if $\mathcal{Q}_{i}$ is balanced and All columns of $\mathcal{Q}_{i}$ non-trivial then
8:              Save $\mathcal{Q}_{i}$.
9:         end if
10:     end for
11:     Keep only unique $\{\mathcal{Q}_{i}\}$.
12:end procedure
3.2 Revisiting Main Problem 1
Let us revisit the problem 1. Given a convex combination of subregion entropies on the LHS with arbitrary positive integer coefficients,
$c_{1}S_{L_{1}}+\cdots+c_{l}S_{L_{l}}.$ (37)
how do we find all possible convex combination of entropies on the RHS $d_{v}\geq 0$ for $\forall v=\{1,\cdots,r\}$,
$d_{1}S_{R_{1}}+\cdots+d_{r}S_{R_{r}},$ (38)
such that
$c_{1}S_{L_{1}}+\cdots+c_{l}S_{L_{l}}\geq d_{1}S_{R_{1}}+\cdots+d_{r}S_{R_{r}}$ (39)
is a valid HEI?
We choose the appropriate hypercube $H_{M}$ corresponding to the LHS and generate all possible contraction maps =$\mathcal{F}=\{f_{\tau=(\sigma,\iota_{N})}\}$. For each contraction map $f_{\tau}$, we assign the same boundary conditions for occurrence vectors, faithfully representing the LHS 37, then each $f_{\tau}$ gives a valid inequality $\mathcal{Q}^{0}_{\tau}$ of the form 39, where the fixed superscript-00 refers to the fixed boundary conditions.
3.3 Example: Deriving the MMI from the contraction map of a star graph
Consider the star graph shown in figure 2. We can construct this graph starting from a hypercube $H_{3}$ and performing the graph contraction with a partition, e.g.,
$\{000\},\{011\},\{101\},\{110\},\{001,010,100,111\}.$ (40)
The resultant star graph is isometrically embeddable in a hypercube $H_{4}$. This embedding can be encoded in the form of a contraction map from $\{0,1\}^{3}$ to $\{0,1\}^{4}$ by (34), given in table 1. The labels $\{L_{u}\}$ and $\{R_{v}\}$ on the contraction map are the LHS and RHS terms of an inequality respectively, to be determined by assigning boundary conditions in $\{s_{k}\}$. We set our convention to assign $s_{1}$ to be the purifier $O$.
$L_{1}$ $L_{2}$ $L_{3}$ $R_{1}$ $R_{2}$ $R_{3}$ $R_{4}$
$s_{1}$ 0 0 0 0 0 0 0
$s_{2}$ 0 0 1 0 0 0 1
$s_{3}$ 0 1 0 0 0 0 1
$s_{4}$ 0 1 1 0 0 1 1
$s_{5}$ 1 0 0 0 0 0 1
$s_{6}$ 1 0 1 1 0 0 1
$s_{7}$ 1 1 0 0 1 0 1
$s_{8}$ 1 1 1 0 0 0 1
Table 1: A contraction map corresponding the star graph shown in figure 2.
We are interested in balanced inequalities that do not contain any trivial $L_{u}$ (and/or $R_{v}$) and thus assign the boundary conditions accordingly. For the choice of the boundary condition $(s_{4},s_{6},s_{7})=(A,B,C)$ and its permutations, we get the MMI inequality,
$S(AB)+S(AC)+S(BC)\geq S(A)+S(B)+S(C)+S(ABC).$ (41)
In fact, for this contraction map (table 1), imposing non-triviality of $\{L_{u}\}$ and $\{R_{v}\}$ yields the MMI inequality as the only candidate, as expected from the uniqueness of the MMI contraction map. We also tabulate some other inequalities derived from the MMI map that with non-trivial column labels in table 2. See Appendix A for further discussions on relaxing the non-triviality of column labels.
Parties List of HEIs
$n=4$
$\{A,B,C,D\}$
$S_{BD}+S_{CD}+S_{ABC}\geq S_{B}+S_{C}+S_{D}+S_{ABCD},$
$S_{BC}+S_{CD}+S_{ABD}\geq S_{B}+S_{C}+S_{D}+S_{ABCD},$
$S_{AC}+S_{AD}+S_{BCD}\geq S_{A}+S_{C}+S_{D}+S_{ABCD}.$
$n=5$
$\{A,B,C,D,E\}$
$S_{DE}+S_{ACD}+S_{BCE}\geq S_{C}+S_{D}+S_{E}+S_{ABCDE},$
$S_{BE}+S_{ABD}+S_{CDE}\geq S_{B}+S_{D}+S_{E}+S_{ABCDE},$
$S_{BD}+S_{ABE}+S_{CDE}\geq S_{B}+S_{D}+S_{E}+S_{ABCDE}.$
$n=6$
$\{A,B,C,D,E,F\}$
$S_{ACE}+S_{BCF}+S_{DEF}\geq S_{C}+S_{E}+S_{F}+S_{ABCDEF}.$
Table 2: The holographic inequalities (up to permutations) for $n=4,5,6$ with $M=3$ and $N=4$, generated from the MMI contraction map (table 1) using algorithm 2.
4 Discussions
4.1 Complexity
We discuss the complexity involved for generating all possible inequalities starting from a hypercube $H_{M}$.
  • There exists a polynomial algorithmer1988 to efficiently generate partitions of a set of cardinality $k$ with computational complexity $\mathcal{O}(k^{1.6})$. Since there are $2^{M}$ vertices in $H_{M}$, the computational complexity associated with the step Partition generator is $\mathcal{O}(2^{M^{1.6}})$
  • The number of all possible partitions is upper bounded by the Bell number, $B_{2^{M}}$. This is an overestimate as multiple contractions can be geometrically equivalent up to rotations of the RHS hypercube. So, the computational complexity for the step Graph contraction is $\mathcal{O}\left(2^{M}\left(B_{2^{M}}\right)\right)$. The asymptotic approximation for the logarithm of Bell numbers of order smaller than $k$ is given by Flajolet_Sedgewick_2009 ,
    $\ln B_{k}=k(\ln k-\ln\ln k-1+o(1)).$ (42)
    where $\ln$ is the natural logarithm.
  • The complexity of finding an embedding (if exists) for a contracted graph goes as $\mathcal{O}(|V|^{2})$ where $|V|$ is the number of verticesEppstein_2011 . Since the number of vertices is upper bounded by $2^{M}$, the step Partial cube identifier has a complexity upper bounded by $\mathcal{O}\left(\left(2^{M}\right)^{2}\right)$. This is an overestimate as graph contractions always reduce the number of vertices.
  • As discussed in section 3.1, the computational complexity of generating possible inequalities from a given inequality is upper bounded by $\mathcal{O}\left((M+N)\left(2^{M}\right)^{n}\right)$. As proved in A.1, we have $M+1\leq N\leq 2^{M-1}$. For facet inequalities, it is an empirical observation that $N<<2^{M-1}$. In principle, one may choose the number of parties $n$ as large as up to $2^{M-1}$, but one is usually interested in tight inequalities, which has empirically been observed to favor small $n$.
Thus the total computational complexity to generate all possible $n$-party HEIs starting from a LHS consisting $M$ terms is upper bounded by $\mathcal{O}\left(B_{2^{M}}\left(2^{M}\right)^{n+3}(M+N)\right)$, where $N$ is the number of RHS terms for some inequality.
However, the total complexity is overestimated, although the algorithm constructs all possible contraction maps as discussed in subsection 2.4. We believe this can be improved with more efficient algorithms, particularly for identifying geometrically equivalent graph contractions, and further results for bounding $n$ for tight inequalities. We leave this for future work.
4.2 Classifications of image graphs $\Phi(H_{M})$
It is a tantalizing direction to study these image graphs and classify them into families of HEIs, shedding further light on their qualitative nature. For example, let us look at the graphs of the contraction maps for the family of cyclic inequalities (see figure 3). All these graphs have a similarity of symmetric structures, standing on one-leg, due to the fact that the entanglement entropy of all labeled regions appears on the RHS. In the $k=3$ case, we have three 1-dimensional edges spreading out, which gets uplifted to five 2-dimensional petals in $k=5$ case, followed by seven 3-dimensional boxes for $k=7$. Predictably, we have a symmetric arrangement of nine 4-dimensional polytopes standing on one-leg for $k=9$. This strategy appears generalizable; whenever there is a family of image graphs $G_{i}$ that can be inductively generated and are isometrically hypercube embeddable, it should correspond to a family of HEIs. We leave this exciting exploration of image graphs and their relation to families of polytopes for future work.
(a) $k=3$
(b) $k=5$
(c) $k=7$
Figure 3: The graphs associated with the first three members of cyclic inequalities.
4.3 Reformulating holographic entropy cones?
The inequalities we generate in this work are more appropriately characterized by the number of unimodular entropies that appear on their LHS’s, rather than their party number. While the full HEC based on fixing party is in principle recoverable in this way, it is somewhat cumbersome to do so. As such, it may be worth reformulating the HEC into one based on the number of unimodular entropies that appear, as opposed to party number, which this algorithmic approach more directly generates.
4.4 Quantum entropy inequalities from contraction maps of HEIs
Recall from section 3.1, that we imposed the non-triviality of columns in our algorithm, so as to avoid running into the ambiguity of interpreting a trivial column. But what happens when we invert this condition? It leads us to potential candidates for quantum inequalities that are valid, not just for holographic states, but all quantum states. First, consider the following proposition about holographic inequalities (see proposition A.1 for proof).
Proposition 4.1 (Lower bound on $N$).
For a facet HEI (except SA) with $M$ terms on the LHS, the number of RHS terms $N$ is bounded below by $M+1\leq N$.
One direct implication is that all such $n$-party facet inequalities (except SA) are violated by the $(n+1)$-party GHZ states (see corollary A.1 for proof). Since, a quantum inequality must always satisfy the GHZ state, these holographic inequalities cannot be valid candidates for quantum inequalities. Thus, we derive a necessary constraint for quantum inequalities.
Corollary 4.1.
For a quantum inequality with $M$ terms on the LHS, the number of RHS terms $N$ is bounded above by $M\geq N$.
At the expense of not interpreting the trivial columns, one can impose this new constraint from corollary 4.1 as a means to generate HEIs with $M\geq N$. These HEIs serve as potential candidates for quantum inequalities. For example, one can recover the subadditivity(1), strong subadditivity(3), Araki-Lieb(2) and weak monotonicity(4) inequalities from the MMI contraction map, all of which are quantum inequalities. This recovery may be explained by the fact that the graph corresponding to these inequalities can be obtained by graph contractions in the MMI graph. We leave a detailed discussion about generating valid quantum inequalities from contraction maps for future work. See Appendix A for an exemplary set of conjectures.
4.5 What could we learn from algorithmic “flatness”?
Consider a set =$\mathcal{F}=\{f|f:\{0,1\}^{M}\to\{0,1\}^{N}\}$ of contraction maps our algorithms generate for fixed $M$. Given a set of disjoint boundary subregions $[n+1]$ including a purifier $O$, we can construct a constant time slice of bulk manifold $\mathcal{M}_{L}$ with $M$ RT surfaces by giving a boundary condition to the contraction map $f\in\mathcal{F}$. Let us fix the boundary conditions for all contraction maps. Then, we can find a set of bulk manifolds $\mathcal{M}_{R}(f)$, with at most $N$ distinct RT surfaces, for each contraction map $f$. The total area of RT surfaces in $\mathcal{M}_{L}$ upper bounds that of RT surfaces in $\mathcal{M}_{R}(f)$ for any $f\in\mathcal{F}$. It thus implies that the number $\eta(\mathcal{M}_{L})$ of different ways to deform the RT surfaces in $\mathcal{M}_{L}$ into those in $\mathcal{M}_{R}(f)$ is at most the number of all the contraction maps, $|\mathcal{F}|$.
Consider two bulk manifolds $\mathcal{M}_{L}$ and $\mathcal{M}_{L}^{\prime}$, which have distinctive configurations of RT surfaces. It could be said that the bulk manifold $\mathcal{M}_{L}$ is more “flat” [^15] In machine learning, flatness of a loss surface characterizes the change in loss under the perturbations of parametersShalev-Shwartz_Ben-David_2014 . than $\mathcal{M}_{L}^{\prime}$ if there are more ways to deform the RT surfaces in $\mathcal{M}_{L}$ than those in $\mathcal{M}_{L}^{\prime}$, i.e., $\eta(\mathcal{M}_{L})\geq\eta(\mathcal{M}_{L}^{\prime})$.
We leave the investigations on what the “flatness” could imply about properties of bulk geometry or vice versa.
4.6 Further speeding up convex optimization
It has previously been shown in Bao:2015bfa that all HEIs are also obeyed by cut functions of general graphs. Therefore, the new classes of inequalities we explore in this work will also be respected by cut functions on graphs. We note that submodular convex optimization is known to be significantly faster than standard convex optimization methods, via the imposition of only the SSA inequality. It is worth investigating whether incorporating these more restrictive inequalities would provide further speed-up to convex optimization approaches to graph theoretic problems.
Acknowledgement
We would like to thank Scott Aaronson, Bartek Czech, Sergio Hernandez-Cuenca, Cynthia Keeler, and Michael Walter for helpful comments. We thank Bartłomiej Czech for comments on the draft. N.B. is supported by DOE ASCR, in particular under the grant Novel Quantum Algorithms from Fast Classical Transforms. K.F. is supported by N.B.’s startup funding at Northeastern University. J.N. is partially supported by the NSF under Cooperative Agreement PHY2019786 and N.B.’s startup funding at Northeastern University.
Appendix A From HEIs to quantum entropy inequalities
In this section, we will discuss quantum entropy inequalities. First we will prove some of our claims from the text.
Definition A.1 (Tripartite form Hernandez-Cuenca:2023iqh ).
An information quantity $\mathcal{Q}$ is said to be in the tripartite form if it is expressed as
$\mathcal{Q}=\sum_{i}-I_{3}(X_{i}:Y_{i}:Z_{i}|W_{i})$ (43)
where the arguments $X_{i},Y_{i},Z_{i},Wi\subset[N]$ are disjoint subsystems, the sum runs over any finite number of terms, and we allow for the conditioning to trivialize, $W_{i}=\emptyset$, in which case $I_{3}(X_{i}:Y_{i}:Z_{i}|\emptyset)=I_{3}(X_{i}:Y_{i}:Z_{i})$ and, they are defined to be
$I_{3}(X_{i}:Y_{i}:Z_{i}|W_{i})=I_{3}(X_{i}:Y_{i}:Z_{i}W_{i})-I_{3}(X_{i}:Y_{i}:W_{i})$ (44)
and,
$I_{3}(X_{i}:Y_{i}:Z_{i})=X_{i}+Y_{i}+Z_{i}-X_{i}Y_{i}-X_{i}Z_{i}-Y_{i}Z_{i}+X_{i}Y_{i}Z_{i}$ (45)
We denote $I^{p}C^{q}$ for a $\mathcal{Q}$ that has $p$ number of $-I_{3}(X_{i}:Y_{i}:Z_{i})$ and $q$ number of $-I_{3}(X_{i}:Y_{i}:Z_{i}|W_{i})$ terms in the sum (43).
We borrow the following conjecture A.1 from Hernandez-Cuenca:2023iqh .
Conjecture A.1.
All facet inequalities (except SA) are expressible in the $I^{p}C^{q}$ form with $p\geq 1$ and $q\geq 0$.
Proposition A.1.
For facet inequalities (except SA), $M+1\leq N$
Proof.
According to conjecture A.1, we can write an inequality $\mathcal{Q}$ of the form $I^{p}C^{q}$ as
$\mathcal{Q}=\sum_{i=1}^{p}-I_{3}(X_{i}:Y_{i}:Z_{i})+\sum_{j=1}^{q}-I_{3}(X_{i}:Y_{i}:Z_{i}|W_{i}):=\sum_{i=1}^{p}I^{(i)}+\sum_{j=1}^{q}C^{(j)}\geq 0,$ (46)
with non-trivial $W_{i}$. It is simple to show that every $C^{(j)}$ has an equal number of positive and negative terms, whereas every $I^{(i)}$ has one more negative term than positive terms, where $i$ and $j$ simply labels the associated $I$ and $C$ terms respectively. Since, positive (negative) terms contribute to LHS (RHS) and $p\geq 1$, we have $M+1\leq N$. ∎
Corollary A.1.
All $n$-party HEIs that are facets of the HEC (except SA) are violated by the $(n+1)$-party GHZ state.
Proof.
For a GHZ, state $S_{X}=S>0$ for all $X\in P(n)\backslash\emptyset$. Since, $N\geq M+1$ for all facet inequalities (except SA), they are trivially violated. ∎
Recall the contraction map of the MMI inequality (given in table 3 below). In this section, we use this map as an example and allow the columns to carry trivial labels such that they satisfy the necessary conditions [^16] We leave the understanding of sufficient conditions for valid quantum entropy inequalities for future work. for quantum entropy inequalities, namely, the existence of a contraction map and constraints on the relative number of terms appearing on two sides,
$L_{1}$ $L_{2}$ $L_{3}$ $R_{1}$ $R_{2}$ $R_{3}$ $R_{4}$
$s_{1}$ 0 0 0 0 0 0 0
$s_{2}$ 0 0 1 0 0 0 1
$s_{3}$ 0 1 0 0 0 0 1
$s_{4}$ 0 1 1 0 0 1 1
$s_{5}$ 1 0 0 0 0 0 1
$s_{6}$ 1 0 1 1 0 0 1
$s_{7}$ 1 1 0 0 1 0 1
$s_{8}$ 1 1 1 0 0 0 1
Table 3: A contraction map corresponding the star graph shown in figure 4(a).
In particular, we will give an example showing how one can arrive at strong subadditivity starting from the contraction map (table 3) corresponding to the MMI. We begin with the MMI, where one may have the the boundary conditions assigned as $(s_{1},s_{4},s_{6},s_{7})=(O,A,B,C)$. For the SSA, the boundary conditions may be changed to $(s_{1},s_{4},s_{6},s_{5})=(O,A,B,C)$ (see figure 4 for a graphical interpretation). The resultant inequality is
$S(AB)+S(BC)\geq S(ABC).$ (47)
(a) MMI
(b) SSA
Figure 4: The star graphs corresponding to MMI (a) and SSA(b) with boundary conditions. One may start with (a), identify the vertices $A\leftrightarrow C$ and contract the edge with the central vertex of the star to get (b).
Similarly, we can also get the subadditivity, weak monotonicity and Araki-Lieb inequalities with the appropriate assignment of boundary conditions. We tabulate some examples of $n=4,5,6$ as potential candidates for quantum inequalities in table 4 generated using a contraction map of the $5$-party cyclic inequality.
Parties Examples of conjectures for quantum inequalities
$n=4$
$\{A,B,C,D\}$
$S_{AB}+S_{BD}+S_{CD}\geq S_{B}+S_{D}+S_{ABCD}$
$S_{BD}+S_{CD}+S_{ABC}\geq S_{D}+S_{BC}+S_{ABCD}$
$n=5$
$\{A,B,C,D,E\}$
$S_{BE}+S_{ABD}+S_{CDE}\geq S_{E}+S_{BD}+S_{ABCDE}$
$S_{AB}+S_{BD}+S_{CE}+S_{DE}\geq S_{B}+S_{D}+S_{E}+S_{ABCDE}$
$n=6$
$\{A,B,C,D,E,F\}$
$S_{ACE}+S_{BCF}+S_{DEF}\geq S_{F}+S_{CE}+S_{ABCDEF}$
$S_{DF}+S_{EF}+S_{ACD}+S_{BCE}\geq S_{E}+S_{F}+S_{CD}+S_{ABCDEF}$
Table 4: Examples of conjectures for quantum inequalities for $n=4,5,6$ generated using a contraction map for the $5$-party cyclic inequality.